OpenStudy (anonymous):

Find the slope and y intercept of -2y=6(5-3x)

OpenStudy (anonymous):

y = mx + b start by distributing the 6

OpenStudy (anonymous):

I did and I got -2y=30-18x

OpenStudy (anonymous):

divide by -2

OpenStudy (anonymous):

-2y=6(5-3x) -2y=30-18x y=-15+9x y=9x-15 m=9, b=y-int=15

OpenStudy (anonymous):

I DID and I got y=30-18x

OpenStudy (anonymous):

u need to divide the 30 and -18x by -2 so y = -15 + 9x y = 9x - 15 what is m?

OpenStudy (anonymous):

Ya thanks I got the answer

OpenStudy (anonymous):

what did u get for the intercept -15?